The lines 3x+2y-1=0 , 4x+5y+3=0 intersect at M find the coordinates of M . Equation of the line through M with 3/4 as it's slope
Answer by Alan3354(69443)

Same problem, different numbers.
********************************************************
The lines 3x+2y-1=0 , 4x+5y=0 intersect at M find the coordinates of M . Equation of the line through M with 3/4 as it's slope
--------------
3x+2y = 1
4x+5y = 0 --> y = -4x/5
===========
3x+2y = 1
sub for y
3x - 8x/5 = 1
7x/5 = 1
x = 5/7
y = -4/7
M is (5/7,-4/7)
=====================
M with 3/4 as it's slope
y - y1 = m*(x - x1)
y + 4/7 = (3/4)*(x - 5/7)

===================
it's = it is